# PRISE: Learning Temporal Action Abstractions as a Sequence Compression Problem

### v1

@@ 5.4 Ablation Analyses
Vocabulary Size. The vocabulary size of the BPE tokenizer, $|\mathcal{V}|$, plays an important role in PRISE. A larger vocabulary size captures more motion patterns but increases the complexity of learning the skill-token policy due to more prediction classes. In addition, it learns less frequent tokens, which may not generalize or useful to downstream tasks. On the other hand, a smaller vocabulary size limits the range of motion patterns that can be captured, but may reduce overfitting (see Section 3.4). In the previous experiments, we set $|\mathcal{V}|$ to be 200. Here, we evaluate five hyperparameter settings $\{50,100,150,200,250,300\}$ for $|\mathcal{V}|$ to assess their impact on downstream policy adaptation, in terms of the averaged 5-shot IL performance across 8 unseen LIBERO tasks. As Figure 8 (left) illustrates, PRISE’s performance remains relatively stable across different hyperparameter settings, except when $|\mathcal{V}|$ is extremely large or small. This demonstrates the robustness of PRISE to the choice of $|\mathcal{V}|$, indicating that as long as the vocabulary size is not too large or small, the performance should not vary significantly.
@@ 6 Conclusion and Future Work
In this work, we propose a new temporal action method for continuous control scenarios,  PRISE, that leverages powerful NLP methodology during pretraining for sequential decision making tasks. By first pretraining a vector quantization module to discretize action into codes and then apply Byte Pair Encoding tokenization algorithm to learn temporally extended skill tokens, PRISE’s pretrained skill tokens can capture diverse motion patterns shared across pretraining tasks, enabling efficient multitask policy learning and few-shot adaptation to unseen tasks. One exciting future direction is to further scale up this approach to large real robot dataset with diverse robot embodiments such as Open Embodiment X (Collaboration et al., 2023). Additionally, instead of finetuning the model to different downstream tasks tabula rasa, we could leverage the pretrained tokens to instruct finetune an existing large language model, so that we could leverage the power of foundational models for generalizing across different tasks and scenarios.
7 Acknowledgements
Zheng and Huang are supported by National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
@@ Appendix A Detailed Related Work
Why learn temporal action abstractions? Learning temporally extended action primitives has a long history in sequential decision-making literature. In fully observable and partially observable MDPs (Puterman, 1994; Sutton & Barto, 2018), the standard mathematical model for decision-making, every action takes one time step to execute. In most RL, IL, and planning algorithms, the decision space consists of individual actions, so standard methods produce policies that make a separate decision at every time step. It has long been informally observed and shown formally under various assumptions that the difficulty of policy learning scales with the problem horizon (Ross et al., 2011; Jiang et al., 2015; Rajaraman et al., 2020; Xie et al., 2021). This motivates hierarchical approaches (Parr & Russell, 1998; Barto & Mahadevan, 2003; Le et al., 2018; Nachum et al., 2018; Kipf et al., 2019; Kujanpää et al., 2023), which view the decision-making process as consisting of a high-level policy that identifies task substeps and invokes lower-level temporally extended routines sometimes called options (Sutton et al., 1999) or primitives (Ajay et al., 2021) to complete them. Conceptually, our PRISE method can be viewed as hierarchical imitation learning (HIL).
@@
Pretraining data assumptions. In terms of the assumptions PRISE imposes on the trajectories it uses for pretraining, it is more similar to Play-LMP (Lynch et al., 2019), RPL (Gupta et al., 2019), MCIL (Lynch & Sermanet, 2021), and TACO-RL (Rosete-Beas et al., 2022) rather than to IL or offline RL. Namely, PRISE needs this data to contain meaningful behavioral patters, which it extracts using an imitation-like procedure (as does RPL). While PRISE could be applied to arbitrary-quality data commonly assumed by offline RL or some state representation pretraining algorithms such as  Schwarzer et al. (2021); Wei et al. (2023); Zheng et al. (2023, 2024), many of the patterns there are unlikely to be useful, and may not easily aggregate into common temporally extended sequences that BPE aims to extract. On the other hand, the pretraining data doesn’t need to be imitation-quality either.
@@ Appendix B Additional Experimental Results
Unseen Tasks
BC (Scratch)
BC (Pretrained)
ACT (Scratch)
ACT (Pretrained)
PRISE w/o BPE
PRISE
Hand Insert
$18.8\pm 12.5$
$46.3\pm 7.6$
$45.6\pm 6.1$
$56.3\pm 2.8$
$54.2\pm 3.9$
$\bm{60.0\pm 5.0}$
Box Close
$19.4\pm 8.7$
$53.9\pm 4.2$
$55.6\pm 6.7$
$58.1\pm 6.8$
$54.8\pm 5.1$
$\bm{60.8\pm 6.6}$
Disassemble
$8.1\pm 17.3$
$65.6\pm 3.6$
$65.2\pm 2.9$
$51.8\pm 8.5$
$69.6\pm 3.2$
$\bm{74.1\pm 7.3}$
Stick Pull
$20.6\pm 5.8$
$56.3\pm 5.6$
$51.8\pm 4.3$
$65.0\pm 4.5$
$59.3\pm 6.2$
$\bm{67.5\pm 5.6}$
Pick Place Wall
$19.4\pm 6.4$
$67.5\pm 3.7$
$56.9\pm 5.9$
$65.2\pm 6.5$
$62.5\pm 4.8$
$\bm{71.7\pm 5.7}$
Mean
$17.3\pm 4.2$
$57.9\pm 1.4$
$55.0\pm 1.8$
$59.3\pm 2.0$
$60.1\pm 1.5$
$\bm{66.8\pm 1.8}$
Table 1: 5-shot imitation learning performance on five unseen tasks in MetaWorld. Results are averaged across 3 random seeds.
@@
Unseen Tasks
BC (Scratch)
BC (Pretrained)
ACT (Scratch)
ACT (Pretrained)
PRISE w/o BPE
PRISE
Task 0
$16.7\pm 6.2$
$6.7\pm 2.3$
$21.7\pm 8.4$
$20.0\pm 6.0$
$16.7\pm 6.2$
$\bm{26.7\pm 6.4}$
Task 1
$25.0\pm 4.7$
$48.3\pm 10.3$
$26.7\pm 10.5$
$33.3\pm 13.1$
$25.0\pm 4.7$
$\bm{48.3\pm 9.4}$
Task 2
$68.3\pm 7.4$
$60.0\pm 4.1$
$61.7\pm 9.2$
$67.7\pm 6.2$
$68.3\pm 6.2$
$\bm{70.0\pm 0.0}$
Task 3
$66.7\pm 7.9$
$66.7\pm 8.4$
$73.3\pm 9.3$
$70.3\pm 6.2$
$66.7\pm 5.9$
$\bm{78.3\pm 8.8}$
Task 4
$25.0\pm 10.8$
$26.7\pm 3.1$
$45.0\pm 4.3$
$35.0\pm 4.1$
$\bm{45.0\pm 10.8}$
$28.3\pm 6.2$
Task 5
$70.0\pm 0.0$
$46.7\pm 13.1$
$58.3\pm 4.3$
$68.3\pm 6.5$
$75.0\pm 4.2$
$\bm{90.0\pm 4.1}$
Task 6
$23.3\pm 4.3$
$21.7\pm 2.4$
$\bm{35.0\pm 4.3}$
$15.0\pm 0.0$
$23.3\pm 6.2$
$25.0\pm 4.1$
Task 7
$28.3\pm 4.3$
$35.0\pm 7.1$
$10.0\pm 0.0$
$26.7\pm 7.0$
$28.3\pm 10.2$
$\bm{45.0\pm 8.1}$
Mean
$40.4\pm 1.2$
$39.0\pm 4.1$
$41.5\pm 2.1$
$42.0\pm 3.2$
$43.5\pm 0.6$
$\bm{52.7\pm 0.9}$
Table 3: 5-shot imitation learning performance across 8 unseen tasks in LIBERO. Resullts are averaged across 3 random seeds
@@ Appendix C Implementation Details
10# beta:             Weight of decoder loss (1.0 in MetaWorld and 0.01 in LIBERO for       #                   numerical stability)
@@
Metaworld: We generate 100 expert trajectories for each pretraining task using the scripted policy provided in MetaWorld. The observation for each step includes an $84\times 84$ third-person view image as well as an 8-dimensional proprioceptive state vector of the robot’s end-effector. We use the same shallow CNN as in Yarats et al. (2022) to encode the observations into a 64-dimensional latent vector and apply a linear layer to embed the 8-dimensional state vector also into a 64-dimensional latent vector. The architecture of the shallow CNN is presented in Figure 12.
@@
LIBERO: For LIBERO, we pretrain PRISE on the provided LIBERO-90 dataset. The pretraining dataset contains 50 demonstration trajectories for each task, collected by human teleoperation. For each timestep, the agent observes a third-person view image, a first-person view image from its wrist camera (both with resolution $128\times 128$), a 9-dimensional vector of the robot’s proprioceptive state, and a task instruction in natural language. We use the exact same architecture as ResNet-T in (Liu et al., 2023). In particular, we use two ResNet18 encoders He et al. (2016) to process the image observations, utilizing FiLM Perez et al. (2018) encoding to incorporate the 768-dimensional BERT embedding of the task language instruction. The action decoder $\psi$ is also a three-layer MLP with hidden size being 1024.
@@ Appendix D Additional Results on Multitask Learning
Task ID
BC (ResNet-RNN)
BC (ResNet-T)
BC (ViT-T)
ACT
PRISE
0
$0.45$
$0.45$
$0.6$
$1.0$
$0.8$
1
$0.1$
$0.1$
$0.0$
$0.6$
$0.35$
2
$0.0$
$0.25$
$0.05$
$0.95$
$0.7$
3
$0.15$
$0.0$
$0.0$
$0.4$
$0.5$
4
$0.2$
$0.0$
$0.0$
$0.3$
$0.45$
5
$0.25$
$0.0$
$0.0$
$0.7$
$0.65$
6
$0.05$
$0.0$
$0.15$
$0.4$
$0.5$
7
$0.3$
$0.45$
$0.25$
$1.0$
$0.95$
8
$0.05$
$0.0$
$0.0$
$1.0$
$0.6$
9
$0.35$
$0.3$
$0.0$
$0.95$
$0.35$
10
$0.2$
$0.7$
$0.15$
$1.0$
$0.95$
11
$0.45$
$0.4$
$0.65$
$1.0$
$0.95$
12
$0.2$
$0.05$
$0.35$
$0.35$
$0.2$
13
$0.1$
$0.35$
$0.0$
$0.25$
$0.4$
14
$0.35$
$0.1$
$0.0$
$0.75$
$0.35$
15
$0.5$
$0.1$
$0.6$
$0.95$
$0.75$
16
$0.25$
$0.05$
$0.25$
$0.75$
$0.4$
17
$0.05$
$0.05$
$0.0$
$0.5$
$0.15$
18
$0.4$
$0.3$
$0.35$
$0.25$
$0.3$
19
$0.15$
$0.0$
$0.4$
$1.0$
$0.65$
20
$0.6$
$0.7$
$1.0$
$1.0$
$1.0$
21
$0.15$
$0.0$
$0.35$
$0.7$
$0.3$
22
$0.75$
$0.4$
$0.55$
$0.75$
$0.85$
23
$0.2$
$0.0$
$0.05$
$0.45$
$0.05$
24
$0.25$
$0.3$
$0.25$
$0.1$
$0.95$
25
$0.8$
$0.6$
$0.85$
$0.1$
$0.9$
26
$0.0$
$0.0$
$0.35$
$0.6$
$0.55$
27
$0.05$
$0.0$
$0.2$
$0.35$
$0.05$
28
$0.05$
$0.8$
$0.9$
$1.0$
$1.0$
29
$0.2$
$0.1$
$0.15$
$0.85$
$1.0$
30
$0.05$
$0.05$
$0.1$
$0.4$
$0.5$
31
$0.5$
$0.25$
$0.2$
$1.0$
$0.85$
32
$0.0$
$0.0$
$0.05$
$0.3$
$0.2$
33
$0.1$
$0.1$
$0.2$
$0.5$
$0.3$
34
$0.2$
$0.05$
$0.1$
$0.5$
$0.8$
35
$0.7$
$0.1$
$0.45$
$1.0$
$0.75$
36
$0.05$
$0.0$
$0.05$
$0.05$
$0.25$
37
$0.05$
$0.05$
$0.15$
$0.0$
$0.3$
38
$0.15$
$0.0$
$0.1$
$0.9$
$0.2$
39
$0.2$
$0.25$
$0.75$
$0.4$
$0.85$
40
$0.45$
$0.15$
$0.05$
$0.9$
$0.5$
41
$0.25$
$0.4$
$0.3$
$0.85$
$0.55$
42
$0.15$
$0.45$
$0.15$
$0.7$
$0.8$
43
$0.1$
$0.1$
$0.15$
$0.85$
$0.4$
44
$0.5$
$0.4$
$0.95$
$0.75$
$0.85$
@@
Task ID
BC (ResNet-RNN)
BC (ResNet-T)
BC (ViT-T)
ACT
PRISE
45
$0.0$
$0.0$
$0.0$
$0.8$
$0.55$
46
$0.0$
$0.0$
$0.0$
$0.0$
$0.35$
47
$0.0$
$0.0$
$0.0$
$0.0$
$0.25$
48
$0.0$
$0.0$
$0.0$
$0.0$
$0.65$
49
$0.0$
$0.0$
$0.0$
$0.0$
$0.65$
50
$0.0$
$0.0$
$0.0$
$0.35$
$0.4$
51
$0.0$
$0.0$
$0.0$
$0.1$
$0.1$
52
$0.0$
$0.0$
$0.0$
$0.05$
$0.3$
53
$0.0$
$0.0$
$0.0$
$0.05$
$0.6$
54
$0.0$
$0.05$
$0.0$
$0.15$
$0.5$
55
$0.0$
$0.05$
$0.0$
$0.0$
$0.35$
56
$0.0$
$0.3$
$0.0$
$0.0$
$0.8$
57
$0.0$
$0.25$
$0.0$
$0.0$
$0.5$
58
$0.0$
$0.0$
$0.0$
$0.5$
$0.2$
59
$0.0$
$0.25$
$0.0$
$0.0$
$0.65$
60
$0.0$
$0.4$
$0.2$
$0.45$
$0.8$
61
$0.0$
$0.2$
$0.25$
$0.05$
$0.85$
62
$0.0$
$0.0$
$0.1$
$0.05$
$0.4$
63
$0.0$
$0.0$
$0.0$
$0.0$
$0.4$
64
$0.0$
$0.0$
$0.05$
$0.25$
$0.15$
65
$0.0$
$0.0$
$0.05$
$0.05$
$0.15$
66
$0.15$
$0.0$
$0.0$
$0.45$
$0.3$
67
$0.1$
$0.15$
$0.05$
$0.55$
$0.55$
68
$0.35$
$0.05$
$0.05$
$0.6$
$0.85$
69
$0.1$
$0.65$
$0.0$
$0.85$
$0.9$
70
$0.05$
$0.55$
$0.0$
$0.6$
$0.55$
71
$0.0$
$0.2$
$0.0$
$0.0$
$0.35$
72
$0.0$
$0.2$
$0.0$
$0.30$
$0.6$
73
$0.0$
$0.0$
$0.05$
$0.35$
$0.3$
74
$0.2$
$0.05$
$0.0$
$0.7$
$0.45$
75
$0.0$
$0.1$
$0.0$
$0.35$
$0.25$
76
$0.0$
$0.3$
$0.35$
$0.1$
$0.65$
77
$0.35$
$0.3$
$0.0$
$0.7$
$0.8$
78
$0.25$
$0.1$
$0.0$
$0.1$
$0.45$
79
$0.25$
$0.45$
$0.05$
$0.95$
$0.3$
80
$0.0$
$0.2$
$0.0$
$0.45$
$0.3$
81
$0.2$
$0.0$
$0.0$
$0.5$
$0.35$
82
$0.4$
$0.45$
$0.15$
$0.55$
$0.8$
83
$0.0$
$0.05$
$0.05$
$0.0$
$0.55$
84
$0.0$
$0.2$
$0.05$
$0.15$
$0.75$
85
$0.0$
$0.0$
$0.05$
$0.1$
$0.75$
86
$0.05$
$0.2$
$0.05$
$0.3$
$0.95$
87
$0.2$
$0.1$
$0.15$
$1.0$
$0.65$
88
$0.0$
$0.25$
$0.05$
$0.85$
$0.55$
89
$0.15$
$0.1$
$0.05$
$0.45$
$0.55$

### v3

@@ 5.4 Ablation Analyses
Vocabulary Size. The vocabulary size of the BPE tokenizer, $|\mathcal{V}|$, plays an important role in PRISE. A larger vocabulary size captures more motion patterns but increases the complexity of learning the skill-token policy due to more prediction classes. In addition, it learns less frequent tokens, which may not generalize or useful to downstream tasks. On the other hand, a smaller vocabulary size limits the range of motion patterns that can be captured, but may reduce overfitting (see Section 3.4). In the previous experiments, we set $|\mathcal{V}|$ to be 200. Here, we evaluate five hyperparameter settings $\{100,150,200,250,300\}$ for $|\mathcal{V}|$ to assess their impact on downstream policy adaptation, in terms of the averaged 5-shot IL performance across 8 unseen LIBERO tasks. As Figure 9 (left) illustrates, PRISE’s performance remains relatively stable across different hyperparameter settings, except when $|\mathcal{V}|$ is extremely large or small. This demonstrates the robustness of PRISE to the choice of $|\mathcal{V}|$, indicating that as long as the vocabulary size is not too large or small, the performance should not vary significantly.
@@
Effects of Dataset Size. To further study how much data is needed to learn a good temporal action abstraction in PRISE, we conduct a set of additional experiments on LIBERO. Instead of using the entire LIBERO-90 dataset, where each task includes 50 human demonstration trajectories, we now sample $N$ trajectories per task, with $N\in\{10,20,30,40,50\}$. We then evaluate the downstream performance of the PRISE skill token. The result is shown in Figure 12.
Figure 12: Downstream performance of PRISE under different pretraining dataset sizes
From the Figure, it is evident that the performance of PRISE monotonically gets better as the number of pretraining trajectories per task increases. Notably, the performance of PRISE is good with up to 40 trajectories per task but drops significantly with fewer than 30 trajectories per task. It would be interesting for future work to further scale up the existing LIBERO-90 dataset so that PRISE may further benefit from an even larger pretraining dataset. But this should be beyond the scope of the current work.
@@ 6 Conclusion and Future Work
In this work, we propose a new temporal action method for continuous control scenarios,  PRISE, that leverages powerful NLP methodology during pretraining for sequential decision making tasks. By first pretraining a vector quantization module to discretize the action space into codes and then applying Byte Pair Encoding to learn temporally extended skill tokens, PRISE captures diverse motion patterns across pretraining tasks, enabling efficient multitask policy learning and few-shot adaptation to unseen problems. One exciting future direction is to further scale up this approach to large real-robot datasets with diverse embodiments, such as Open X-Embodiment (Padalkar et al., 2024). Additionally, instead of finetuning the model to different downstream tasks tabula rasa, we could leverage the pretrained tokens to instruction-finetune an existing large language model, so that we could capitalize on its generalization power across different tasks and scenarios.
Impact Statement
This paper presents a pretraining algorithm to advance policy learning for sequential decision making problems. The proposed method is demonstrated in experiments of simulation robotic manipulation. As it is, this work does not have significant societal consequences. However, when the method here is scaled up to train large, foundation models, it would build AI/robotics systems that can more efficiently adapt and generalize. Such an improvement may pose potential societal risks (such as taking over human labour).
Acknowledgements
Ruijie Zheng and Furong Huang are supported by National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships.
@@ Appendix A Detailed Related Work
Why learn temporal action abstractions? Learning temporally extended action primitives has a long history in sequential decision-making literature. In fully observable and partially observable MDPs (Puterman, 1994; Sutton & Barto, 2018), the standard mathematical model for decision-making, every action takes one time step to execute. In most RL, IL, and planning algorithms, the decision space consists of individual actions, so standard methods produce policies that make a separate decision at every time step. It has long been informally observed and shown formally under various assumptions that the difficulty of policy learning scales with the problem horizon (Ross et al., 2011; Jiang et al., 2015; Rajaraman et al., 2020; Xie et al., 2021). This motivates hierarchical approaches (Parr & Russell, 1998; Barto & Mahadevan, 2003; Le et al., 2018; Nachum et al., 2018; Kipf et al., 2019; Kujanpää et al., 2023; Jiang et al., 2022), which view the decision-making process as consisting of a high-level policy that identifies task substeps and invokes lower-level temporally extended routines sometimes called options (Sutton et al., 1999) or primitives (Ajay et al., 2021) to complete them. Conceptually, our PRISE method can be viewed as hierarchical imitation learning (HIL).
@@
Pretraining data assumptions. In terms of the assumptions PRISE imposes on the trajectories it uses for pretraining, it is more similar to C-BeT (Cui et al., 2023), Play-LMP (Lynch et al., 2019), RPL (Gupta et al., 2019), MCIL (Lynch & Sermanet, 2021), and TACO-RL (Rosete-Beas et al., 2022) rather than to IL or offline RL. Namely, PRISE needs this data to contain meaningful behavioral patters, which it extracts using an imitation-like procedure (as does RPL). While PRISE could be applied to arbitrary-quality data commonly assumed by offline RL or some state representation pretraining algorithms such as  (Schwarzer et al., 2021; Sun et al., 2023; Wei et al., 2023; Zheng et al., 2023a, 2024; Hansen et al., 2022b; Yuan et al., 2022), many of the patterns there are unlikely to be useful, and may not easily aggregate into common temporally extended sequences that BPE aims to extract. On the other hand, the pretraining data doesn’t need to be imitation-quality either.
@@ Appendix B Additional Experimental Results
Unseen Tasks BC (Scratch) BC (Pretrained) ACT (Scratch) ACT (Pretrained) PRISE w/o BPE PRISE
Hand Insert $18.8\pm 12.5$ $46.3\pm 7.6$ $45.6\pm 6.1$ $56.3\pm 2.8$ $54.2\pm 3.9$ $\bm{60.0\pm 5.0}$
Box Close $19.4\pm 8.7$ $53.9\pm 4.2$ $55.6\pm 6.7$ $58.1\pm 6.8$ $54.8\pm 5.1$ $\bm{60.8\pm 6.6}$
Disassemble $8.1\pm 17.3$ $65.6\pm 3.6$ $65.2\pm 2.9$ $51.8\pm 8.5$ $69.6\pm 3.2$ $\bm{74.1\pm 7.3}$
Stick Pull $20.6\pm 5.8$ $56.3\pm 5.6$ $51.8\pm 4.3$ $65.0\pm 4.5$ $59.3\pm 6.2$ $\bm{67.5\pm 5.6}$
Pick Place Wall $19.4\pm 6.4$ $67.5\pm 3.7$ $56.9\pm 5.9$ $65.2\pm 6.5$ $62.5\pm 4.8$ $\bm{71.7\pm 5.7}$
Mean $17.3\pm 4.2$ $57.9\pm 1.4$ $55.0\pm 1.8$ $59.3\pm 2.0$ $60.1\pm 1.5$ $\bm{66.8\pm 1.8}$
Table 1: 5-shot imitation learning performance on five unseen tasks in MetaWorld. Results are averaged across 4 random seeds.
@@
Unseen Tasks BC (Scratch) BC (Pretrained) ACT (Scratch) ACT (Pretrained) PRISE w/o BPE PRISE
Task 0 $16.7\pm 6.2$ $6.7\pm 2.3$ $21.7\pm 8.4$ $20.0\pm 6.0$ $16.7\pm 6.2$ $\bm{26.7\pm 6.4}$
Task 1 $25.0\pm 4.7$ $48.3\pm 10.3$ $26.7\pm 10.5$ $33.3\pm 13.1$ $25.0\pm 4.7$ $\bm{48.3\pm 9.4}$
Task 2 $68.3\pm 7.4$ $60.0\pm 4.1$ $61.7\pm 9.2$ $67.7\pm 6.2$ $68.3\pm 6.2$ $\bm{70.0\pm 0.0}$
Task 3 $66.7\pm 7.9$ $66.7\pm 8.4$ $73.3\pm 9.3$ $70.3\pm 6.2$ $66.7\pm 5.9$ $\bm{78.3\pm 8.8}$
Task 4 $25.0\pm 10.8$ $26.7\pm 3.1$ $45.0\pm 4.3$ $35.0\pm 4.1$ $\bm{45.0\pm 10.8}$ $28.3\pm 6.2$
Task 5 $70.0\pm 0.0$ $46.7\pm 13.1$ $58.3\pm 4.3$ $68.3\pm 6.5$ $75.0\pm 4.2$ $\bm{90.0\pm 4.1}$
Task 6 $23.3\pm 4.3$ $21.7\pm 2.4$ $\bm{35.0\pm 4.3}$ $15.0\pm 0.0$ $23.3\pm 6.2$ $25.0\pm 4.1$
Task 7 $28.3\pm 4.3$ $35.0\pm 7.1$ $10.0\pm 0.0$ $26.7\pm 7.0$ $28.3\pm 10.2$ $\bm{45.0\pm 8.1}$
Mean $40.4\pm 1.2$ $39.0\pm 4.1$ $41.5\pm 2.1$ $42.0\pm 3.2$ $43.5\pm 0.6$ $\bm{52.7\pm 0.9}$
Table 3: 5-shot imitation learning performance across 8 unseen tasks in LIBERO. Resullts are averaged across 4 random seeds
@@ Appendix C Implementation Details
10# beta: Weight of decoder loss (1.0 in MetaWorld and 0.01 in LIBERO for numerical stability)
@@
Note: Here our decision to adapt a BYOL-like objective for the forward dynamics losses, rather than directly modeling the forward transition dynamics as in model-based learning approaches (Janner et al., 2019; Zheng et al., 2023b; Hansen et al., 2022a; Hafner et al., 2023; Wang et al., 2024), is mainly motivated by the effectiveness of the BYOL (Schwarzer et al., 2021) objective in preventing representation collapse during latent forward dynamics training, as we are jointly learning both the action quantization and state representation here in the pretraining stage I.
Metaworld: We generate 100 expert trajectories for each pretraining task using the scripted policy provided in MetaWorld. The observation for each step includes an $84\times 84$ third-person view image as well as an 8-dimensional proprioceptive state vector of the robot’s end-effector. We use the same shallow CNN as in (Yarats et al., 2022) to encode the observations into a 64-dimensional latent vector and apply a linear layer to embed the 8-dimensional state vector also into a 64-dimensional latent vector. The architecture of the shallow CNN is presented in Figure 14.
@@
LIBERO: For LIBERO, we pretrain PRISE on the provided LIBERO-90 dataset. The pretraining dataset contains 50 demonstration trajectories for each task, collected by human teleoperation. For each timestep, the agent observes a third-person view image, a first-person view image from its wrist camera (both with resolution $128\times 128$), a 9-dimensional vector of the robot’s proprioceptive state, and a task instruction in natural language. We use the exact same architecture as ResNet-T in (Liu et al., 2023). In particular, we use two ResNet18 encoders (He et al., 2016) to process the image observations, utilizing FiLM (Perez et al., 2018) encoding to incorporate the 768-dimensional BERT embedding of the task language instruction. The action decoder $\psi$ is also a three-layer MLP with hidden size being 1024.
@@ Appendix D Additional Results on Multitask Learning
Task ID BC (ResNet-RNN) BC (ResNet-T) BC (ViT-T) ACT PRISE
0 $0.45$ $0.45$ $0.6$ $1.0$ $0.8$
1 $0.1$ $0.1$ $0.0$ $0.6$ $0.35$
2 $0.0$ $0.25$ $0.05$ $0.95$ $0.7$
3 $0.15$ $0.0$ $0.0$ $0.4$ $0.5$
4 $0.2$ $0.0$ $0.0$ $0.3$ $0.45$
5 $0.25$ $0.0$ $0.0$ $0.7$ $0.65$
6 $0.05$ $0.0$ $0.15$ $0.4$ $0.5$
7 $0.3$ $0.45$ $0.25$ $1.0$ $0.95$
8 $0.05$ $0.0$ $0.0$ $1.0$ $0.6$
9 $0.35$ $0.3$ $0.0$ $0.95$ $0.35$
10 $0.2$ $0.7$ $0.15$ $1.0$ $0.95$
11 $0.45$ $0.4$ $0.65$ $1.0$ $0.95$
12 $0.2$ $0.05$ $0.35$ $0.35$ $0.2$
13 $0.1$ $0.35$ $0.0$ $0.25$ $0.4$
14 $0.35$ $0.1$ $0.0$ $0.75$ $0.35$
15 $0.5$ $0.1$ $0.6$ $0.95$ $0.75$
16 $0.25$ $0.05$ $0.25$ $0.75$ $0.4$
17 $0.05$ $0.05$ $0.0$ $0.5$ $0.15$
18 $0.4$ $0.3$ $0.35$ $0.25$ $0.3$
19 $0.15$ $0.0$ $0.4$ $1.0$ $0.65$
20 $0.6$ $0.7$ $1.0$ $1.0$ $1.0$
21 $0.15$ $0.0$ $0.35$ $0.7$ $0.3$
22 $0.75$ $0.4$ $0.55$ $0.75$ $0.85$
23 $0.2$ $0.0$ $0.05$ $0.45$ $0.05$
24 $0.25$ $0.3$ $0.25$ $0.1$ $0.95$
25 $0.8$ $0.6$ $0.85$ $0.1$ $0.9$
26 $0.0$ $0.0$ $0.35$ $0.6$ $0.55$
27 $0.05$ $0.0$ $0.2$ $0.35$ $0.05$
28 $0.05$ $0.8$ $0.9$ $1.0$ $1.0$
29 $0.2$ $0.1$ $0.15$ $0.85$ $1.0$
30 $0.05$ $0.05$ $0.1$ $0.4$ $0.5$
31 $0.5$ $0.25$ $0.2$ $1.0$ $0.85$
32 $0.0$ $0.0$ $0.05$ $0.3$ $0.2$
33 $0.1$ $0.1$ $0.2$ $0.5$ $0.3$
34 $0.2$ $0.05$ $0.1$ $0.5$ $0.8$
35 $0.7$ $0.1$ $0.45$ $1.0$ $0.75$
36 $0.05$ $0.0$ $0.05$ $0.05$ $0.25$
37 $0.05$ $0.05$ $0.15$ $0.0$ $0.3$
38 $0.15$ $0.0$ $0.1$ $0.9$ $0.2$
39 $0.2$ $0.25$ $0.75$ $0.4$ $0.85$
40 $0.45$ $0.15$ $0.05$ $0.9$ $0.5$
41 $0.25$ $0.4$ $0.3$ $0.85$ $0.55$
42 $0.15$ $0.45$ $0.15$ $0.7$ $0.8$
43 $0.1$ $0.1$ $0.15$ $0.85$ $0.4$
44 $0.5$ $0.4$ $0.95$ $0.75$ $0.85$
@@
Task ID BC (ResNet-RNN) BC (ResNet-T) BC (ViT-T) ACT PRISE
45 $0.0$ $0.0$ $0.0$ $0.8$ $0.55$
46 $0.0$ $0.0$ $0.0$ $0.0$ $0.35$
47 $0.0$ $0.0$ $0.0$ $0.0$ $0.25$
48 $0.0$ $0.0$ $0.0$ $0.0$ $0.65$
49 $0.0$ $0.0$ $0.0$ $0.0$ $0.65$
50 $0.0$ $0.0$ $0.0$ $0.35$ $0.4$
51 $0.0$ $0.0$ $0.0$ $0.1$ $0.1$
52 $0.0$ $0.0$ $0.0$ $0.05$ $0.3$
53 $0.0$ $0.0$ $0.0$ $0.05$ $0.6$
54 $0.0$ $0.05$ $0.0$ $0.15$ $0.5$
55 $0.0$ $0.05$ $0.0$ $0.0$ $0.35$
56 $0.0$ $0.3$ $0.0$ $0.0$ $0.8$
57 $0.0$ $0.25$ $0.0$ $0.0$ $0.5$
58 $0.0$ $0.0$ $0.0$ $0.5$ $0.2$
59 $0.0$ $0.25$ $0.0$ $0.0$ $0.65$
60 $0.0$ $0.4$ $0.2$ $0.45$ $0.8$
61 $0.0$ $0.2$ $0.25$ $0.05$ $0.85$
62 $0.0$ $0.0$ $0.1$ $0.05$ $0.4$
63 $0.0$ $0.0$ $0.0$ $0.0$ $0.4$
64 $0.0$ $0.0$ $0.05$ $0.25$ $0.15$
65 $0.0$ $0.0$ $0.05$ $0.05$ $0.15$
66 $0.15$ $0.0$ $0.0$ $0.45$ $0.3$
67 $0.1$ $0.15$ $0.05$ $0.55$ $0.55$
68 $0.35$ $0.05$ $0.05$ $0.6$ $0.85$
69 $0.1$ $0.65$ $0.0$ $0.85$ $0.9$
70 $0.05$ $0.55$ $0.0$ $0.6$ $0.55$
71 $0.0$ $0.2$ $0.0$ $0.0$ $0.35$
72 $0.0$ $0.2$ $0.0$ $0.30$ $0.6$
73 $0.0$ $0.0$ $0.05$ $0.35$ $0.3$
74 $0.2$ $0.05$ $0.0$ $0.7$ $0.45$
75 $0.0$ $0.1$ $0.0$ $0.35$ $0.25$
76 $0.0$ $0.3$ $0.35$ $0.1$ $0.65$
77 $0.35$ $0.3$ $0.0$ $0.7$ $0.8$
78 $0.25$ $0.1$ $0.0$ $0.1$ $0.45$
79 $0.25$ $0.45$ $0.05$ $0.95$ $0.3$
80 $0.0$ $0.2$ $0.0$ $0.45$ $0.3$
81 $0.2$ $0.0$ $0.0$ $0.5$ $0.35$
82 $0.4$ $0.45$ $0.15$ $0.55$ $0.8$
83 $0.0$ $0.05$ $0.05$ $0.0$ $0.55$
84 $0.0$ $0.2$ $0.05$ $0.15$ $0.75$
85 $0.0$ $0.0$ $0.05$ $0.1$ $0.75$
86 $0.05$ $0.2$ $0.05$ $0.3$ $0.95$
87 $0.2$ $0.1$ $0.15$ $1.0$ $0.65$
88 $0.0$ $0.25$ $0.05$ $0.85$ $0.55$
89 $0.15$ $0.1$ $0.05$ $0.45$ $0.55$